# DressCode: Autoregressively Sewing and Generating Garments from Text Guidance

### v2

@@ Abstract.
Apparel’s significant role in human appearance underscores the importance of garment digitalization for digital human creation. Recent advances in 3D content creation are pivotal for digital human creation. Nonetheless, garment generation from text guidance is still nascent. We introduce a text-driven 3D garment generation framework, DressCode, which aims to democratize design for novices and offer immense potential in fashion design, virtual try-on, and digital human creation. For our framework, we first introduce SewingGPT, a GPT-based architecture integrating cross-attention with text-conditioned embedding to generate sewing patterns with text guidance. We also tailored a pre-trained Stable Diffusion for high-quality, tile-based PBR texture generation. By leveraging a large language model, our framework generates CG-friendly garments through natural language interaction. Our method also facilitates pattern completion and texture editing, streamlining the design process through user-friendly interaction. This framework fosters innovation by allowing creators to freely experiment with designs and incorporate unique elements into their work, thereby igniting new ideas and artistic possibilities. With comprehensive evaluations and comparisons with other state-of-the-art methods, our method showcases the best quality and alignment with input prompts. User studies further validate our high-quality rendering results, highlighting its practical utility and potential in production settings. Our project page is https://IHe-KaiI.github.io/DressCode/.
@@
copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4publicationmonth: 7doi: 10.1145/3658147
@@ 1. Introduction
In contrast, for our graphics community, the dominant representation of garments is sewing patterns, which facilitates both physical simulation and animation in a CG-friendly fashion (Blender Foundation, 2022; Autodesk, INC., 2019). For sewing pattern generation, early methods (Umetani et al., 2011; Berthouzoz et al., 2013) only use simple partial modules in the workflow like parsing or draping to 3D. With the consolidation of more advanced datasets (Korosteleva and Lee, 2021), very recent work enables sewing pattern generation from point clouds (Korosteleva and Lee, 2022) or images (Liu et al., 2023d). (Korosteleva and Sorkine-Hornung, 2023) design a DSL for sewing pattern construction, allowing parameterization and algorithmic support. However, they largely overlook the generation through more natural language interactions, let alone handling the vivid generation with desired texture patterns or physically-based materials, which could significantly speed up the preliminary stage of garment design. Furthermore, novices without professional skills in design or complex design software need to describe their ideas only in text prompts to transform these ideas into designs. This significantly lowers the design barriers, allowing more people to participate in the creative process. Most importantly, generative models introduce diversity to designed text prompts, generate varied types of garments conforming to the prompts, and stimulate designer creativity.
@@
To achieve high-quality garment rendering, we progressively tailor a pre-trained Stable Diffusion model (Rombach et al., 2022) to generate tile-based physically-based rendering (PBR) textures from text prompts. We first fine-tune the U-Net architecture (Ronneberger et al., 2015) of the diffusion model within latent space to generate the diffuse attribute, and then fine-tune the various VAE (Kingma and Welling, 2013) decoders to generate normal and roughness maps separately.
We showcase the capability of DressCode to generate CG-friendly garments with rich sewing patterns and PBR textures from text prompts. We also demonstrate the versatility of our approach, including a ChatGPT-like conversational agent for interactive garment generation, garment completion from partial inputs, and user-friendly texture editing. To summarize, our main contributions include:
@@ 4. Customized Garment Generation
With SewingGPT, we have the capability to generate diverse sewing patterns directly from text prompts. Recognizing appearance’s crucial role in the CG pipeline, we aim to generate corresponding physically-based textures for these patterns, aligning more closely with garment design workflows. By leveraging the SewingGPT and PBR texture generator, our framework DressCode further utilizes a large language model to create customized garments for users through natural language interaction.
@@ 5. Experiments
In this section, we first conduct qualitative and quantitative comparison experiments with other state-of-the-art 3D generation methods to demonstrate the generation capability of our method. We then present ablation studies to evaluate our pipeline. Furthermore, we conduct a comprehensive user study to showcase our results compared to other methods. We also verify our results in Marvelous Designer (CLO3D, 2024); we manually load generated garments into the software and then animate the garments with the human model in non-T-pose. Finally, we show a results gallery in Figure 9 of generated high-quality garments using our method with various text prompts, including generated sewing patterns, PBR textures, draping results on a T-pose human model, and animated results with various poses under different illuminations.
@@ 5.5. User Study
We conduct a comprehensive user study to evaluate the quality of our generated customized garments, particularly their alignment with given text prompts and overall quality. Given 20 text prompts, including descriptions of the shape and texture of garments, we then render the generated results for each method and shuffle the order of results obtained by different methods, as shown in Figure 15. Then we ask approximately 30 users to select the best results from all candidates with comprehensive consideration for two aspects: conformity to the captions of prompt texts regarding garments’ shape and texture, and the visualized quality and fidelity of the rendered garments. As illustrated in Table 2, the preference results clearly indicate a significant advantage of our method over competing approaches in both aspects, highlighting its superiority in aligning closely with the textual prompts and producing visually appealing and high-fidelity garments.
@@ Acknowledgements.
This work was supported by the National Key R&D Program of China (2022YFF0902301), NSFC programs (61976138, 61977047), STCSM (2015F0203-000-06), and SHMEC(2019-01-07-00-01-E00003). We also acknowledge support from the Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI).

### v4

@@ Abstract.
Apparel’s significant role in human appearance underscores the importance of garment digitalization for digital human creation. Recent advances in 3D content creation are pivotal for digital human creation. Nonetheless, garment generation from text guidance is still nascent. We introduce a text-driven 3D garment generation framework, DressCode, which aims to democratize design for novices and offer immense potential in fashion design, virtual try-on, and digital human creation. We first introduce SewingGPT, a GPT-based architecture integrating cross-attention with text-conditioned embedding to generate sewing patterns with text guidance. We then tailor a pre-trained Stable Diffusion to generate tile-based Physically-based Rendering (PBR) textures for the garments. By leveraging a large language model, our framework generates CG-friendly garments through natural language interaction. It also facilitates pattern completion and texture editing, streamlining the design process through user-friendly interaction. This framework fosters innovation by allowing creators to freely experiment with designs and incorporate unique elements into their work. With comprehensive evaluations and comparisons with other state-of-the-art methods, our method showcases superior quality and alignment with input prompts. User studies further validate our high-quality rendering results, highlighting its practical utility and potential in production settings. Our project page is https://IHe-KaiI.github.io/DressCode/.
@@
copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 72publicationmonth: 7doi: 10.1145/3658147ccs: Computing methodologies Computer graphics
@@ 1. Introduction
In contrast, for our graphics community, the dominant representation of garments is sewing patterns, which facilitates both physical simulation and animation in a CG-friendly fashion (Blender Foundation, 2022; Autodesk, INC., 2019). For sewing pattern generation, early methods (Umetani et al., 2011; Berthouzoz et al., 2013) only use simple partial modules in the workflow like parsing or draping to 3D. With the consolidation of more advanced datasets (Korosteleva and Lee, 2021), recent work enables sewing pattern generation from point clouds (Korosteleva and Lee, 2022) or images (Liu et al., 2023d). However, they largely overlook the generation through more natural language interactions, let alone handling the vivid generation with desired texture patterns or physically-based materials, which could significantly speed up the preliminary stage of garment design. Furthermore, through the natural interactions of text prompts, novices without professional skills in complex design software can directly describe and transform their ideas into creations. This significantly lowers the design barriers, allowing more newcomers to participate in the creative process. Most importantly, generative models introduce diversity to designed text prompts, generate varied types of garments conforming to the prompts, and hence stimulate designer creativity.
@@
To achieve high-quality garment rendering, we progressively tailor a pre-trained Stable Diffusion model (Rombach et al., 2022) to generate tile-based Physically-based Rendering (PBR) textures from text prompts. We first fine-tune the U-Net architecture (Ronneberger et al., 2015) of the diffusion model within latent space to generate the diffuse attribute, and then fine-tune the various VAE (Kingma and Welling, 2013) decoders to generate normal and roughness maps separately. We showcase the capability of DressCode to generate CG-friendly garments with rich sewing patterns and PBR textures from text prompts. We also demonstrate the versatility of our approach, including a ChatGPT-like conversational agent for interactive garment generation, garment completion from partial inputs, and user-friendly texture editing. To summarize, our main contributions include:
@@ 4. Customized Garment Generation
With SewingGPT, we have the capability to generate diverse sewing patterns directly from text prompts. Recognizing appearance’s crucial role in the CG pipeline, we aim to generate corresponding Physically-based Rendering (PBR) textures for these patterns, aligning more closely with garment design workflows. By leveraging the SewingGPT and PBR texture generator, our framework DressCode further utilizes a large language model to create customized garments for users through natural language interaction.
@@ 5. Experiments
In this section, we first conduct qualitative and quantitative comparison experiments with other state-of-the-art 3D generation methods to demonstrate the generation capability of our method. We then present ablation studies and validation to evaluate our pipeline. Furthermore, we conduct a qualitative comparison experiment with parametric templates and perform a comprehensive user study to showcase our results compared to other methods. We also verify our results in Marvelous Designer (CLO3D, 2024); we manually load generated garments into the software and then animate the garments with the human model in non-T-pose. We then show a results gallery in Figure 9 of generated high-quality garments using our method with various text prompts, including generated sewing patterns, PBR textures, draping results on a T-pose human model, and animated results with various poses under different illuminations.
@@ 5.5. User Study
We conduct a comprehensive user study to evaluate the quality of our generated customized garments, particularly their alignment with given text prompts and overall quality. Given 20 text prompts, including descriptions of the shape and texture of garments, we then render the generated results for each method and shuffle the order of results obtained by different methods, as shown in Figure 15. Then we ask 30 users to select the best results from all candidates with comprehensive consideration for two aspects: conformity to the captions of prompt texts regarding garments’ shape and texture, and the visualized quality and fidelity of the rendered garments. As illustrated in Table 2, the preference results clearly indicate a significant advantage of our method over competing approaches in both aspects, highlighting its superiority in aligning closely with the textual prompts and producing visually appealing and high-fidelity garments.
@@ Acknowledgements.
This work was supported by the National Key R&D Program of China (2022YFF0902301), NSFC programs (61976138, 61977047), STCSM (2015F0203-000-06), and SHMEC(2019-01-07-00-01-E00003). We also acknowledge support from the Shanghai Frontiers Science Center of Human-centered Artificial Intelligence and MoE Key Lab of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University).